OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). y = 2x + 6 y = -1/2x + 3 y = −-1/2x + 6 y = 2x + 3 Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2). y = -1/3x − 6 y = −1/3x − 14 y = 3x − 6 y = 3x − 14

OpenStudy (anonymous):

perpendicular m1 * m2 = -1 Parallel m1 = m2

OpenStudy (anonymous):

Consider y=2x+6 Gradient of the above line = 2 Therefor, gradient of any line perpendicular to the above line = -1/2 Since the considered perpendicular line is passing through (6,3) ; the required line , (y-3) =(-1/2)*(x-6) 2y+x-12=0 You can do the next question similarly

OpenStudy (shamim):

ok i wanna try

OpenStudy (shamim):

for ur question 1

OpenStudy (anonymous):

@shamim I just need the answer to #2 now sorry

OpenStudy (shamim):

ur given equation of line is y=2x+2

OpenStudy (shamim):

ok

OpenStudy (shamim):

ur given line of equation is 3x-y=-3

OpenStudy (shamim):

or,y=3x+3

OpenStudy (shamim):

so the slope of ur line is m=3

OpenStudy (shamim):

and the slope of a parralel line is also m=3

OpenStudy (shamim):

and this parallel line will go through a point (4,-2)

OpenStudy (anonymous):

I don't it to be explained, sorry, I just need the answer.. and I need it QUICK. Sorry

OpenStudy (shamim):

so the equation of ur parralel line will be y-(-2)=3(x-4)

OpenStudy (shamim):

answer will b y=3x-14